# When Less Language is More: Language-Reasoning Disentanglement Makes LLMs Better Multilingual Reasoners

### v1

@@ Appendix A Further Discussion, Limitations, and Future Works
While our focus is on multilingual reasoning tasks, the principle of decoupling task-relevant and language-specific signals may extend to other high-level cognitive abilities in LLMs, such as planning, reflection, or explanation generation. Investigating how this approach generalizes to other domains, and whether other types of entangled information can be removed, offers a promising avenue for broader interpretability-guided model improvement.
@@
Connection to Latent Reasoning
Our findings on language–reasoning decoupling resonate with the emerging paradigm of latent reasoning in LLMs [Deng et al., 2024, Goyal et al., 2024, Shen et al., 2025a, b]. Recent work by Hao et al. [2024] introduces the Coconut framework, which enables LLMs to perform reasoning within a continuous latent space, rather than relying solely on explicit language tokens. This approach allows models to internally process reasoning steps as high-dimensional representations—termed “continuous thoughts”—before generating any output.
This convergence opens up promising avenues for future research, such as integrating language–reasoning decoupling techniques with latent reasoning frameworks to further improve the generalization and interpretability of LLMs across diverse tasks and languages.
@@ Appendix B Broader Impact
Our work addresses this issue by investigating the internal mechanisms behind multilingual reasoning in LLMs and introducing a lightweight, training-free approach to enhance reasoning capabilities across languages. By explicitly decoupling language-specific signals from reasoning processes, we show that reasoning can be made more language-invariant—without requiring costly additional data or retraining. This has the potential to democratize access to strong reasoning models for underrepresented languages, particularly in settings where resources for large-scale finetuning are limited.
@@
Nonetheless, our intervention raises new questions about fairness and control: the choice to remove language-specific features must be carefully balanced against preserving cultural and linguistic identity. Future work should examine the societal implications of language–reasoning decoupling in real-world multilingual applications, and how such techniques may interact with biases in training data or user experience.
@@ Appendix C Probing for Language Subspace
The optimal solution of Equation 2 can be computed efficiently via Singular Value Decomposition (SVD). Algorithm 1 presents the detailed procedure. Readers interested in more details can consult the proof provided in Xie et al. [2022]. The only hyperparameter $r<L$ controls the amount of language-specific information captured by the identified subspace. The larger $r$ is, the more language-specific signals we can identify.
@@ Appendix D Representation Space Visualization
Figure LABEL:fig:rep_analysis_qwen25 shows the visualizations of Qwen-2.5-7B-Instruct at three different layer depths: lower, middle, and upper. Across all layers, we observe a consistent trend—after projection, multilingual representations become less language-specific and tend to converge toward the English cluster. However, the degree of convergence varies: the middle layers exhibit the most prominent language alignment, while the lower layers show weaker separation to begin with, and the upper layers retain stronger language-specific traits even after ablation. These results suggest that while the decoupling effect holds across the model, the middle layers offer the most effective separation between language and reasoning representations.
@@
Figures LABEL:fig:rep_analysis_qwen25_3b, Figures LABEL:fig:rep_analysis_qwen3_8b, Figures LABEL:fig:rep_analysis_r1_7b and LABEL:fig:rep_analysis_qwq present analogous visualizations for R1-Distill-Qwen-7B and QwQ-32B, respectively. The same convergence trend can be observed: representations of various languages tend to collapse toward the English representation after projection, especially in the middle layers. This confirms that the observed phenomenon generalizes beyond a single model family or training paradigm.
@@ Appendix E Benchmark
  • MGSM [Shi et al., 2023]: [^1] https://huggingface.co/datasets/juletxara/mgsm The Multilingual Grade School Math Benchmark (MGSM) is a dataset designed to evaluate the reasoning abilities of large language models in multilingual settings. It comprises 250 grade-school math problems, originally from the GSM8K dataset, each translated by human annotators into 10 diverse languages: English, Spanish, French, German, Russian, Chinese, Japanese, Thai, Swahili, Bengali, and Telugu . These problems require multi-step reasoning, making MGSM a valuable resource for assessing models’ capabilities in mathematical problem-solving across different languages. The dataset includes both the questions and their step-by-step solutions, facilitating comprehensive evaluation of multilingual reasoning performance.
  • XWinograd [Muennighoff et al., 2023]: [^2] https://huggingface.co/datasets/Muennighoff/xwinograd A well-established tool for evaluating coreference resolution (CoR) and commonsense reasoning (CSR) capabilities of computational models. The dataset is the translation of the English Winograd Schema datasets and it adds 488 Chinese schemas from CLUEWSC2020 [Xu et al., 2020], totaling 6 languages. Formulated as a fill-in-a-blank task with binary options, the goal is to choose the right option for a given sentence which requires commonsense reasoning. In our experimental setup, this benchmark covers English (En), French (Fr), Chinese (Zh) and Japanese (Jp) and Russian (Ru).
  • M-MMLU [Hendrycks et al., 2021, Lai et al., 2023]: [^3] https://huggingface.co/datasets/alexandrainst/m_mmlu A benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings. The datasets is a machine translated version of the MMLU dataset by GPT-3.5-turbo and covers 34 languages. This is a massive multitask test consisting of multiple-choice questions from various branches of knowledge. To attain high accuracy on this test, models must possess extensive world knowledge and problem solving ability. In our experimental setup, this benchmark covers English (En), Spanish (Es), French (Fr), German (De), Chinese (Zh), Japanese (Jp), Bengali (Bn) and Swahili (Sw).
@@ Appendix F Implementation Details of Causal Intervention
We use 7,500 math problems from the Google Translate version of the MATH dataset [Hendrycks et al., 2021], translated into 10 languages—Bengali (bn), German (de), Spanish (es), French (fr), Japanese (ja), Russian (ru), Swahili (sw), Telugu (te), Thai (th), and Chinese (zh)—as the data for extracting the language-specific subspace. For each input, we ablate activations only on the tokens in the prompt. In the main experiments, we perform a grid search to find the best results, applying a $\lambda$ in the range of 0 to 0.4 at the middle layers and -0.4 to 0 at the higher layers. Furthermore, in the analysis of ablation strength presented in Section 4.1, our focus is on evaluating model performance in the middle layers and examining language fidelity in the higher layers.
@@ G.1 Correlation Analysis on Each Language
Figure LABEL:fig:strength_analysis_qwq presents the effects of varying ablation strength on both reasoning performance and output fidelity for each of the 10 evaluation languages, using QwQ-32B as the backbone model. Overall, we observe that increasing ablation strength—i.e., removing more language-specific components—leads to improved reasoning performance across most languages, consistent with the aggregate trend discussed in Section 4.1. However, the fidelity curves show more nuanced patterns, particularly across languages of different resource levels.
@@
We further replicate this per-language analysis on two additional models: R1-Distill-Qwen-7B and Qwen-2.5-Instruct-7B, as shown in Figure LABEL:fig:strength_analysis_r1_7b and Figure LABEL:fig:strength_analysis_qwen25_7b, respectively. Across both models, we observe consistent trends with those reported for QwQ-32B: increasing ablation strength generally improves multilingual reasoning performance, while response fidelity declines more significantly in lower-resource languages.
These results reinforce the generality of our findings across different models and training paradigms. They further suggest that the entanglement between language-specific activation and reasoning is a robust phenomenon, and that language–reasoning decoupling can benefit diverse LLMs—though it may require language-aware tuning to preserve fluency in less represented languages.
@@ G.2 Layer-wise Analysis on More Models
To further verify the robustness of our findings, we extend the layer-wise analysis introduced in Section 4.2 to additional models, including Qwen-2.5-3B-Instruct, Qwen-3-1.7B, Qwen-3-8B, and R1-Distill-Qwen-14B. For each model, we apply the same intervention procedure by ablating language-specific components at different layer depths—specifically lower, middle, and upper layers—and evaluate the effects on multilingual reasoning performance and output fidelity. Results are shown in Figure LABEL:fig:layer_analysis_app.
@@ Appendix H Configuration for Post-training
All training experiments are conducted on eight A100 GPUs using the LLaMA-Factory repository [Zheng et al., 2024]. For distributed training, we leverage the DeepSpeed [Rasley et al., 2020] framework with ZeRo-2 optimization. The optimizer is AdamW. We train the model for 3 epochs on the multilingual version of the MATH dataset (7,500 samples), with a total batch size of 32, a learning rate of 1e-5 and the max context length of 8,192. The learning rate follows a cosine annealing schedule with 10% warm-up steps.
@@
For RL, we apply the PPO algorithm [Schulman et al., 2017] using the OpenRLHF framework [Hu et al., 2024]. Each training run consists of 3 episodes, with 4 rollouts per sample. We train for 1 epoch on the same multilingual MATH dataset, using a batch size of 96. The maximum input length is set to 1,024 tokens, and the maximum output length is 8,192 tokens to accommodate long reasoning traces. The actor model is optimized with a learning rate of 5e-7, while the critic uses a higher learning rate of 9e-6. We apply a KL penalty coefficient of 0.01 to stabilize training and prevent the actor from drifting too far from the initial policy.

### v2

@@ Appendix A Further Discussion, Limitations, and Future Work
While our focus is on multilingual reasoning tasks, the principle of disentangling task-relevant and language-specific signals may extend to other high-level cognitive abilities in LLMs, such as planning, reflection, or explanation generation. Investigating how this approach generalizes to other domains, and whether other types of entangled information can be removed, offers a promising avenue for broader interpretability-guided model improvement.
@@ Appendix B Broader Impact
Our work addresses this issue by investigating the internal mechanisms behind multilingual reasoning in LLMs and introducing a lightweight, training-free approach to enhance reasoning capabilities across languages. By explicitly performing language-reasoning disentanglement, we show that reasoning can be made more language-invariant—without requiring costly additional data or retraining. This has the potential to democratize access to strong reasoning models for underrepresented languages, particularly in settings where resources for large-scale finetuning are limited.
@@
Nonetheless, our intervention raises new questions about fairness and control: the choice to remove language-specific features must be carefully balanced against preserving cultural and linguistic identity. Future work should examine the societal implications of language–reasoning disentanglement in real-world multilingual applications, and how such techniques may interact with biases in training data or user experience.
@@ Appendix C Probing for Language Subspace
The optimal solution of Equation 2 can be computed efficiently via Singular Value Decomposition (SVD). Algorithm 1 presents the detailed procedure. Readers interested in more details can consult the proof provided in xie2022discovering. The only hyperparameter $r<L$ controls the amount of language-specific information captured by the identified subspace. The larger $r$ is, the more language-specific signals we can identify.
@@ 1 $\boldsymbol{M}_{a}^{\prime}\leftarrow\frac{1}{d}\boldsymbol{M}\boldsymbol{\mathbbm{1}}$;
(a) Layer-wise PCA visualizations in Qwen-2.5-7B-Instruct. Each subfigure shows hidden states at lower, middle, and upper layers, before and after projection.
@@ Appendix D Representation Space Visualization
Figure 7(a) shows the visualizations of Qwen-2.5-7B-Instruct at three different layer depths: lower, middle, and upper. Across all layers, we observe a consistent trend—after projection, multilingual representations become less language-specific and tend to converge toward the English cluster. However, the degree of convergence varies: the middle layers exhibit the most prominent language alignment, while the lower layers show weaker separation to begin with, and the upper layers retain stronger language-specific traits even after ablation. These results suggest that while the disentanglement effect holds across the model, the middle layers offer the most effective separation between language and reasoning representations.
@@
Figures 7(b), Figures 7(c), Figures 7(d) and 7(e) present analogous visualizations for R1-Distill-Qwen-7B and QwQ-32B, respectively. The same convergence trend can be observed: representations of various languages tend to collapse toward the English representation after projection, especially in the middle layers. This confirms that the observed phenomenon generalizes beyond a single model family or training paradigm.
(b) Layer-wise PCA visualizations in Qwen-2.5-3B-Instruct. Each subfigure shows hidden states at lower, middle, and upper layers, before and after projection.
(c) Layer-wise PCA visualizations in Qwen-3-8B-Thinking. Each subfigure shows hidden states at lower, middle, and upper layers, before and after projection.
(d) Layer-wise PCA visualizations in R1-Distill-Qwen-7B. Each subfigure shows hidden states at lower, middle, and upper layers, before and after projection.
(e) Layer-wise PCA visualizations in QwQ-32B. Each subfigure shows hidden states at lower, middle, and upper layers, before and after projection.
@@ Appendix E Benchmark
  • MGSM [shi2023language]: [^1] https://huggingface.co/datasets/juletxara/mgsm The Multilingual Grade School Math Benchmark (MGSM) is a dataset designed to evaluate the reasoning abilities of large language models in multilingual settings. It comprises 250 grade-school math problems, originally from the GSM8K dataset, each translated by human annotators into 10 diverse languages: English, Spanish, French, German, Russian, Chinese, Japanese, Thai, Swahili, Bengali, and Telugu . These problems require multi-step reasoning, making MGSM a valuable resource for assessing models’ capabilities in mathematical problem-solving across different languages. The dataset includes both the questions and their step-by-step solutions, facilitating comprehensive evaluation of multilingual reasoning performance.
  • XWinograd [muennighoff2023crosslingual]: [^2] https://huggingface.co/datasets/Muennighoff/xwinograd A well-established tool for evaluating coreference resolution (CoR) and commonsense reasoning (CSR) capabilities of computational models. The dataset is the translation of the English Winograd Schema datasets and it adds 488 Chinese schemas from CLUEWSC2020 [xu2020clue], totaling 6 languages. Formulated as a fill-in-a-blank task with binary options, the goal is to choose the right option for a given sentence which requires commonsense reasoning. In our experimental setup, this benchmark covers English (En), French (Fr), Chinese (Zh) and Japanese (Jp) and Russian (Ru).
  • M-MMLU [hendrycks2021measuring, lai2023okapi]: [^3] https://huggingface.co/datasets/alexandrainst/m_mmlu A benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings. The datasets is a machine translated version of the MMLU dataset by GPT-3.5-turbo and covers 34 languages. This is a massive multitask test consisting of multiple-choice questions from various branches of knowledge. To attain high accuracy on this test, models must possess extensive world knowledge and problem solving ability. In our experimental setup, this benchmark covers English (En), Spanish (Es), French (Fr), German (De), Chinese (Zh), Japanese (Jp), Bengali (Bn) and Swahili (Sw).
@@ Appendix F Implementation Details of Causal Intervention
We use 7,500 math problems from the Google Translate version of the MATH dataset [hendrycks2021measuring], translated into 10 languages—Bengali (bn), German (de), Spanish (es), French (fr), Japanese (jp), Russian (ru), Swahili (sw), Telugu (te), Thai (th), and Chinese (zh)—as the data for extracting the language-specific subspace. For each input, we ablate activations only on the tokens in the prompt. In the main experiments, we perform a grid search to find the best results, applying a $\lambda$ in the range of 0 to 0.4 at the middle layers and -0.4 to 0 at the higher layers. For various LLMs, the specific range of middle and high layers we have selected are shown in Table 4. Furthermore, in the analysis of ablation strength presented in Section 4.1, our focus is on evaluating model performance in the middle layers and examining language fidelity in the higher layers. Also, We conduct significance tests on the multilingual results, showing that our method yields statistically significant gains ($p<0.05$) on most benchmarks. We believe this further supports the effectiveness of our approach, especially given its low cost and wide language coverage.
Table 4: The exact layer ranges for each model, detailing the total number of layers along with specified middle and higher layer ranges.
Model Name Number of Layers Middle Layers Higher Layers
Qwen-2.5-Instruct-3B 36 12-26 27-35
Qwen-2.5-Instruct-7B 28 10-19 20-27
Qwen-3-1.7B-Thinking 28 10-19 20-27
Qwen-3-4B-Thinking 36 12-26 27-35
Qwen-3-8B-Thinking 36 12-26 27-35
R1-Distill-Qwen-7B 28 10-19 20-27
R1-Distill-LLama-8B 32 12-22 23-31
R1-Distill-Qwen-14B 48 16-33 34-47
GLM-Z1-9B 40 12-30 31-39
QwQ-32B 64 20-46 47-63
(a) Effects of ablation strength on each language. The backbone is QwQ-32B.
(b) Effects of ablation strength on each language. The backbone is R1-Distill-Qwen-7B.
(c) Effects of ablation strength on each language. The backbone is Qwen-2.5-7B-Instruct.
@@ G.1 Correlation Analysis on Each Language
Figure 16(a) presents the effects of varying ablation strength on both reasoning performance and output fidelity for each of the 10 evaluation languages, using QwQ-32B as the backbone model. Overall, we observe that increasing ablation strength—i.e., removing more language-specific components—leads to improved reasoning performance across most languages, consistent with the aggregate trend discussed in Section 4.1. However, the fidelity curves show more nuanced patterns, particularly across languages of different resource levels.
@@
We further replicate this per-language analysis on two additional models: R1-Distill-Qwen-7B and Qwen-2.5-Instruct-7B, as shown in Figure 16(b) and Figure 16(c), respectively. Across both models, we observe consistent trends with those reported for QwQ-32B: increasing ablation strength generally improves multilingual reasoning performance, while response fidelity declines more significantly in lower-resource languages.
These results reinforce the generality of our findings across different models and training paradigms. They further suggest that the entanglement between language-specific activation and reasoning is a robust phenomenon, and that language–reasoning disentanglement can benefit diverse LLMs—though it may require language-aware tuning to preserve fluency in less represented languages.
(d) Layer-wise impact of language–reasoning disentanglement on MGSM accuracy and output fidelity. The x-axis denotes the starting layer index of the intervention. Most layers support effective disentanglement that improves reasoning performance.
@@ G.2 Layer-wise Analysis on More Models
To further verify the robustness of our findings, we extend the layer-wise analysis introduced in Section 4.2 to additional models, including Qwen-2.5-3B-Instruct, Qwen-3-1.7B, Qwen-3-8B, and R1-Distill-Qwen-14B. For each model, we apply the same intervention procedure by ablating language-specific components at different layer depths—specifically lower, middle, and upper layers—and evaluate the effects on multilingual reasoning performance and output fidelity. Results are shown in Figure 16(d).
@@
G.3 Impact of Dataset Difficulty and Translation Quality
While prior experiments primarily focus on established multilingual reasoning datasets such as MGSM, these benchmarks may not fully capture the difficulty spectrum of modern reasoning tasks. To more comprehensively assess our approach under varying task complexity and translation fidelity, we further evaluate on PolyMath [wang2025polymath], a recently introduced human-verified multilingual dataset for mathematical reasoning.
The PolyMath dataset contains problems annotated with three difficulty levels—low, medium, and high—allowing us to examine reasoning robustness across linguistic and cognitive complexity. For each target language, we randomly sample 50 questions per difficulty level and compute a difficulty-weighted accuracy score. Experiments are conducted on two representative models, DeepSeek-R1-Distill-Qwen-7B and Qwen-2.5-Instruct-7B, both before and after language–reasoning disentanglement.
Results in Table 5 show that higher-quality and difficulty-balanced datasets substantially reduce the reasoning performance gap between English and other high-resource languages (e.g., Spanish, French). However, mid- and low-resource languages such as Swahili and Telugu continue to underperform, suggesting that structural capability disparities persist even with improved data. Importantly, our intervention helps close this gap: after language–reasoning disentanglement, weaker languages (e.g., Russian in DeepSeek, Thai in Qwen) exhibit notable gains, achieving performance levels comparable to stronger languages. These findings demonstrate both the practical value and the robustness of our approach under more controlled and linguistically diverse evaluation conditions.
Table 5: Multilingual reasoning performance on PolyMath datasets across different languages, before and after language-reasoning disentanglement within the activation spaces of the backbone models (+ L-R Disentangle). The best results are highlighted in bold. The values in parentheses indicate language fidelity to indicate input-output consistency.
High-Resource Mid-Resource Low-Resource AVG.
En Es Fr De Zh Jp Ru Th Te Bn Sw -
Qwen-2.5-Instruct-7B 23.14 23.14 22.29 21.43 18.29 20.00 28.86 16.57 14.29 17.14 9.43 19.51 (42.00%)
+ L-R Disentangle 25.71 28.86 25.14 20.57 22.00 23.43 28.00 21.43 17.14 18.57 16.29 22.47 (42.00%)
R1-Distill-Qwen-7B 48.86 46.29 41.14 49.14 45.14 45.71 39.71 42.29 28.29 42.29 21.14 40.91 (41.94%)
+ L-R Disentangle 50.29 50.86 48.00 51.14 46.57 43.43 46.86 41.71 34.00 45.14 21.71 43.31 (44.18%)
@@ Appendix H Configuration for Post-training
All training experiments are conducted on eight A100 GPUs using the LLaMA-Factory repository [zheng2024llamafactory]. For distributed training, we leverage the DeepSpeed [rasley2020deepspeed] framework with ZeRo-2 optimization. The optimizer is AdamW. We train the model for 3 epochs on the multilingual version of the MATH dataset (7,500 samples), with a total batch size of 32, a learning rate of 1e-5 and the max context length of 8,192. The learning rate follows a cosine annealing schedule with 10% warm-up steps.
@@
For RL, we apply the PPO algorithm [schulman2017proximal] using the OpenRLHF framework [hu2024openrlhf]. Each training run consists of 3 episodes, with 4 rollouts per sample. We train for 1 epoch on the same multilingual MATH dataset, using a batch size of 96. The maximum input length is set to 1,024 tokens, and the maximum output length is 8,192 tokens to accommodate long reasoning traces. The actor model is optimized with a learning rate of 5e-7, while the critic uses a higher learning rate of 9e-6. We apply a KL penalty coefficient of 0.01 to stabilize training and prevent the actor from drifting too far from the initial policy.